7.The Holtzman Corporation has assets of $418,000, current liabilities of $126,000, and long-term liabilities of $131,000. There is $38,700 in preferred stock outstanding; 20,000 shares of common stock have been issued.  


a. Compute book value (net worth) per share. (Round your answer to 2 decimal places.)

Book value per share _________?

b. If there is $32,300 in earnings available to common stockholders, and Holtzman’s stock has a P/E of 21 times earnings per share, what is the current price of the stock? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Current Price ________?

c. What is the ratio of market value per share to book value per share? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Market value to book value _______ times?
Answer #1

a) Book value per share S 6.12 =(418000-126000-131000-38700)/20000 b) Current Price $33.92 =(32300/20000)*21 c) Market value

*Please rate thumbs up